### v1

# The Skolem Problem in rings of positive characteristic

Abstract
We show that the Skolem Problem is decidable in finitely generated commutative rings of positive characteristic. More precisely, we show that there exists an algorithm which, given a finite presentation of a (unitary) commutative ring $\mathcal{R}=\mathbb{Z}_{/T}[X_{1},\ldots,X_{n}]/I$ of characteristic $T>0$, and a linear recurrence sequence $(\gamma_{n})_{n\in\mathbb{N}}\in\mathcal{R}^{\mathbb{N}}$, determines whether $(\gamma_{n})_{n\in\mathbb{N}}$ contains a zero term. Our proof is based on two recent results: Dong and Shafrir (2025) on the solution set of S-unit equations over $p^{e}$-torsion modules, and Karimov, Luca, Nieuwveld, Ouaknine, and Worrell (2025) on solving linear equations over powers of two multiplicatively independent numbers. Our result implies, moreover, that the zero set of a linear recurrence sequence over a ring of characteristic $T=p_{1}^{e_{1}}\cdots p_{k}^{e_{k}}$ is effectively a finite union of $p_{i}$-normal sets in the sense of Derksen (2007).
1 Introduction and main results
The Skolem Problem in a commutative ring $\mathcal{R}$ is the problem of determining whether a given linear recurrence sequence $(\gamma_{n})_{n\in\mathbb{N}}$ over $\mathcal{R}$ contains a zero term. The Skolem Problem is deeply related to central questions in program verification [OW15, AKK+21], control theory [BT00, FOP+19], dynamical systems [KL86, AAGT15], and number theory [Bel06, LMN+23]. Despite its simple formulation, decidability of the Skolem Problem over $\mathbb{Z}$ (and equivalently, $\mathbb{Q}$) is a major open problem in mathematics and computer science: see [Mah35, Lec53, TMS84, BP02, LLN+22, LOW25] for a selected list of partial results and recent progress. Indeed, the celebrated Skolem-Mahler-Lech theorem [Sko34] shows that the zero set of a linear recurrence sequence over $\mathbb{Z}$ is a union of a finite set and finitely many arithmetic progressions. However, the Skolem-Mahler-Lech theorem is not effective, meaning there is currently no known computable bound on this finite set, and hence the decidability of the Skolem Problem over the ring $\mathbb{Z}$ remains open.
In contrast, decidability of the Skolem Problem in fields of prime characteristic (such as $\mathbb{F}_{p}(X)$) was shown by Derksen [Der07], who gave an effective description of the zero set as a $p$-normal set. In this paper we continue Derksen’s work in this direction by showing decidability of the Skolem Problem in rings of arbitrary positive characteristics. In contrast to fields, the characteristic of such a ring is not necessarily prime (e.g., the polynomial ring $\mathbb{Z}_{/6}[X]$ has characteristic $6$). Thus a difficulty is to describe the interaction of different prime divisors of the characteristic.
Let $\mathcal{R}$ be a commutative ring with unity $1$. The characteristic of the ring $\mathcal{R}$ is defined as the smallest positive integer $T$ such that $T\cdot 1=0$ in $\mathcal{R}$. If no such positive integer exists, the ring is said to have characteristic zero. A sequence $\gamma=(\gamma_{0},\gamma_{1},\gamma_{2},\ldots)\in\mathcal{R}^{\mathbb{N}}$ is called an linear recurrence sequence if there exist $d\geq 1$ and $a_{1},\ldots,a_{d}\in\mathcal{R}$ with $a_{d}\neq 0$, such that
$\gamma_{n}=a_{1}\gamma_{n-1}+\cdots+a_{d}\gamma_{n-d}$ (1.1)
for all $n\geq d$. Note that the sequence $\gamma$ is uniquely determined by the recurrence relation (1.1) as well as its $d$ initial terms $\gamma_{0},\ldots,\gamma_{d-1}$. Since every term in $\gamma$ is an element in the subring generated by $a_{1},\ldots,a_{d},\gamma_{0},\ldots,\gamma_{d-1}$, by restricting to this subring we can suppose $\mathcal{R}$ to be finitely generated. The main result of this paper is:
Theorem 1.1.
Let $\mathcal{R}$ be a finitely generated commutative (unitary) ring of characteristic $T>0$. Given a linear recurrence sequence $\gamma\in\mathcal{R}^{\mathbb{N}}$, it is decidable whether $\gamma$ contains a zero.
We say a few words about how the ring $\mathcal{R}$ is represented. Denote by $\mathbb{Z}_{/T}$ the quotient ring $\mathbb{Z}/T\mathbb{Z}=\{0,1,\ldots,T-1\}$ and let $r_{1},\ldots,r_{N}$ be the generators of $\mathcal{R}$. Since $T\cdot 1=0$ in $\mathcal{R}$, we have a surjective homomorphism $\varphi$ from the polynomial ring $\mathbb{Z}_{/T}[X_{1},\ldots,X_{N}]$ to $\mathcal{R}$, defined by $X_{i}\mapsto r_{i},i=1,\ldots,N$. Therefore $\mathcal{R}$ can be written as a quotient $\mathbb{Z}_{/T}[X_{1},\ldots,X_{N}]/I$, with $I=\ker(\varphi)$. This ideal $I$ is finitely generated since $\mathbb{Z}_{/T}[X_{1},\ldots,X_{N}]$ is Noetherian. Therefore, throughout this paper, the ring $\mathcal{R}$ is always represented as a quotient $\mathbb{Z}_{/T}[X_{1},\ldots,X_{N}]/I$, where the generators of $I$ are explicitly given.
Theorem 1.1 for rings of prime characteristics can be deduced from the case of fields following Derksen’s result. By [Der07, Section 9], given any linear recurrence sequence $\gamma$ over a finitely generated ring of prime characteristic $p$, its zero set $\mathfrak{Z}(\gamma)\coloneqq\{n\in\mathbb{N}\mid\gamma_{n}=0\}$ is effectively $p$-normal (see Section 2 for a formal definition). For example, the linear recurrence sequence $\gamma$ over the polynomial ring $\mathbb{F}_{p}[X]$, defined by $\gamma_{n}=(X+1)^{n}-X^{n}-1$, admits the zero set $\{n=p^{k}\mid k\in\mathbb{N}\}$ [Der07, Example 1.3].
Overview of main contributions and proof structure.
More generally, when considering a linear recurrence sequence $\gamma$ over a ring of non-prime characteristic $T=p_{1}^{e_{1}}\cdots p_{k}^{e_{k}}$, we can use the Chinese Remainder Theorem to decompose the zero set $\mathfrak{Z}(\gamma)$ as an intersection $\mathfrak{Z}(\alpha_{1})\cap\cdots\cap\mathfrak{Z}(\alpha_{k})$ of zero sets, where each $\alpha_{i}$ is a linear recurrence sequence over a ring of characteristic $p_{i}^{e_{i}}$. Then, in order to decide whether $\gamma$ contains a zero (i.e., whether $\mathfrak{Z}(\gamma)$ is non-empty), one is immediately confronted with two major obstacles.
The first obstacle is that Derksen’s result does not apply to rings of prime-power characteristics $p^{e},e\geq 2$, due to its reliance on the Frobenius homomorphism. To overcome this, in this paper we will extend Derksen’s result from prime characteristic to prime-power characteristics $p^{e}$. Namely, we will show the following.
Proposition 1.2.
Let $p$ be a prime number, $e$ be a positive integer, and $\mathcal{A}$ be a finitely generated commutative unitary ring of characteristic $p^{e}$. For any given linear recurrence sequence $\alpha\in\mathcal{A}^{\mathbb{N}}$, its zero set $\mathfrak{Z}(\alpha)\coloneqq\{n\in\mathbb{N}\mid\alpha_{n}=0\}$ is effectively $p$-normal.
The formal definition of $p$-normal sets will be given in Section 2 (Definition 2.1). An example of $p$-normal sets is $\{1+5\cdot p^{2a}+p^{2b}\mid a,b\in\mathbb{N}\}$. In particular, these are special forms of $p$-automatic sets [WB00], and is a generalization of linear combinations of powers of $p$. Our key argument to proving Proposition 1.2 is a deep result of Dong and Shafrir [DS25, Theorem 1.3], who showed that the solution set to an S-unit equation over a $p^{e}$-torsion module is effectively $p$-normal.
The second obstacle is to compute the intersection $\mathfrak{Z}(\alpha_{1})\cap\cdots\cap\mathfrak{Z}(\alpha_{k})$ of the zero sets. Due to Proposition 1.2, this boils down to computing the intersection of $p_{i}$-normal sets for different primes $p_{1},\ldots,p_{k}$. In general, it is unknown whether one can effectively compute the intersection of $p_{i}$-automatic sets for different $p_{i}$’s [HS22, AB23]. However, we will show that this is possible in the special case of $p_{i}$-normal sets. In particular, we will prove that every intersection of $p_{i}$-normal sets can actually be written as a union of $p_{i}$-normal sets:
Proposition 1.3.
Let $p_{1},\ldots,p_{k}\in\mathbb{Z}_{>0}$ be multiplicatively independent positive integers. For $i=1,\ldots,k$, let $S_{i}$ be a $p_{i}$-normal subset of $\mathbb{N}$. Then the intersection $S_{1}\cap\cdots\cap S_{k}$ is effectively equal to a union $T_{1}\cup\cdots\cup T_{k}$, where each $T_{i}$ is a $p_{i}$-normal subset of $\mathbb{N}$. In particular, it is decidable whether the intersection $S_{1}\cap\cdots\cap S_{k}$ is empty.
Our main tool for proving Proposition 1.3 comes from a recent paper by Karimov, Luca, Nieuwveld, Ouaknine, and Worrell [KLN+25], which gave an effective procedure to decide the existential fragment of Presburger arithmetic extended with two power predicates. For example, this allows an effective resolution of equations of the form $1+5\cdot p_{1}^{2a}+p_{1}^{2b}=3+p_{2}^{c}$, over the variables $a,b,c\in\mathbb{N}$. However, the effective procedure in [KLN+25] does not extend to more than two power predicates. Therefore, additional insights are needed to compute the intersection of $p_{i}$-normal sets for $k\geq 3$ distinct primes $p_{1},\ldots,p_{k}$.
Combining Proposition 1.2 and 1.3, one also obtains an effective characterization of the zero set of a linear recurrence sequence over a ring $\mathcal{R}$ of characteristic $T=p_{1}^{e_{1}}\cdots p_{k}^{e_{k}}$: namely it is effectively a finite union of $p_{i}$-normal subset of $\mathbb{N}$ for $i=1,\ldots,k$.
2 Preliminaries
2.1 $p$-normal sets
We define $p$-normal subsets of $\mathbb{Z}^{n}$, $\mathbb{Z}$, and $\mathbb{N}$, following the convention established in [Der07] and [DM15].
Definition 2.1 (reformulation of [DM15]).
A set $S\subseteq\mathbb{Z}^{n}$ is called elementary $p$-nested if it is of the form
$\{\boldsymbol{a}_{0}+p^{\ell k_{1}}\boldsymbol{a}_{1}+\cdots+p^{\ell k_{r}}\boldsymbol{a}_{r}\mid k_{1},k_{2},\ldots,k_{r}\in\mathbb{N}\},$ (2.1)
where $\ell\geq 1$ and $\boldsymbol{a}_{0},\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{r}\in\mathbb{Q}^{n}$. A singleton is by definition elementary $p$-nested with $r=0$. Note that the entries of $\boldsymbol{a}_{i}$’s do not have to be integers: for example, the set $\left\{\frac{1}{2}+3^{k}\cdot\frac{1}{2}\;\middle|\;k\in\mathbb{N}\right\}\subseteq\mathbb{Z}$ is elementary $p$-nested.
A subset $S$ of $\mathbb{Z}^{n}$ is called $p$-succinct, if it is of the form $H+D\coloneqq\{\boldsymbol{h}+\boldsymbol{d}\mid\boldsymbol{h}\in H,\boldsymbol{d}\in D\}$, where $H$ is a subgroup of $\mathbb{Z}^{n}$ and $D$ is elementary $p$-nested. For example, the set $\{(x,x+p^{k})\mid x\in\mathbb{Z},k\in\mathbb{N}\}\subseteq\mathbb{Z}^{2}$ is $p$-succinct since it is the sum of the subgroup $H=\{(x,x)\mid x\in\mathbb{Z}\}\leq\mathbb{Z}^{2}$ and the elementary $p$-nested set $D=\{(0,1)\cdot p^{k}\mid k\in\mathbb{N}\}$.
A subset $S$ of $\mathbb{Z}^{n}$ is called $p$-normal if it is a finite union of $p$-succinct sets. The empty set is by definition $p$-normal. A set is called effectively $p$-normal if all its defining coefficients can be effectively computed.
Specializing Definition 2.1 at $n=1$, we obtain a simpler characterization of $p$-normal subsets of $\mathbb{Z}$. Indeed, a $p$-succinct subset $S=H+D\subseteq\mathbb{Z}$ is either elementary $p$-nested (if the subgroup $H$ is trivial), or it is a finite union of cosets $(a\mathbb{Z}+b_{1})\cup\cdots\cup(a\mathbb{Z}+b_{k})$ with $a\geq 1$ (if the subgroup $H$ is of the form $a\mathbb{Z}$). Therefore, $p$-normal subsets of $\mathbb{Z}$ can be equivalently defined as follows.
Observation 2.2.
A subset $S$ of $\mathbb{Z}$ is $p$-normal, if and only if it is a finite union of elementary $p$-nested sets and sets of the form $a\mathbb{Z}+b$, $a\geq 1,b\in\mathbb{Z}$.
For subsets of $\mathbb{N}$, the definition is slightly more technical:
Definition 2.3.
A subset $S$ of $\mathbb{N}$ is called $p$-normal, if there exists $n\geq 0$, a finite set $F\subseteq\{0,1,\ldots,n-1\}$, and a $p$-normal subset $S^{\prime}$ of $\mathbb{Z}$, such that
$S=\left(S^{\prime}\cap[n,\infty)\right)\cup F.$
In other words, a set $S\subseteq\mathbb{N}$ is $p$-normal if it ultimately coincides with a $p$-normal subset of $\mathbb{Z}$. Therefore various closure properties for $p$-normal sets:
Lemma 2.4 ([Der07, Lemma 9.5]).
Let $S_{1},S_{2}$ be $p$-normal subsets of $\mathbb{N}$ (or $\mathbb{Z}$). Then the intersection $S_{1}\cap S_{2}$ is effectively $p$-normal.
Lemma 2.5 (Special case of [DM15, Proposition 2.1]).
Let $\boldsymbol{b}\in\mathbb{Z}^{n}$, $B$ be a subgroup of $\mathbb{Z}^{n}$ and $U\subseteq\mathbb{Z}^{n}$ be an elementary $p$-nested set. Then, the set of integers $\{z\in\mathbb{Z}\mid z\cdot\boldsymbol{b}\in B+U\}$ is effectively $p$-normal.
From Lemma 2.5, we immediately obtain:
Corollary 2.6.
Let $S\subseteq\mathbb{Z}^{n}$ be a $p$-normal set and let $\boldsymbol{b}\in\mathbb{Z}^{n}$. Then, the set of integers $\{z\in\mathbb{Z}\mid z\cdot\boldsymbol{b}\in S\}$ is effectively $p$-normal.
Proof.
Since $p$-normal subsets of $\mathbb{Z}$ are closed under finite union, it suffices to prove the case where $S$ is $p$-succinct, which corresponds exactly to Lemma 2.5. ∎
2.2 commutative algebra
All rings considered in this paper are commutative and unitary. We recall some standard definitions from commutative algebra [Eis13].
Definition 2.7.
Let $R$ be an effective commutative Noetherian ring.
  1. (i)
    An element $r\in R$ is called a zero-divisor if there exists a non-zero $x$ in $R$ such that $rx=0$. For example, $3$ is a zero-divisor of the ring $\mathbb{Z}/12\mathbb{Z}$.
  2. (ii)
    An element $r\in R$ is called nilpotent if there exists $n\in\mathbb{N}$ such that $r^{n}=0$. For example, $6$ is nilpotent in $\mathbb{Z}/12\mathbb{Z}$.
  3. (iii)
    A proper ideal $I\subsetneq R$ is called prime if for every $a,b\in R$ such that $ab\in I$, either $a\in I$ or $b\in I$. Equivalently, $I$ is prime if the ring $R/I$ has no zero-divisor other than zero.
  4. (iv)
    A proper ideal $I\subsetneq R$ is called primary if for every $a,b\in R$ such that $ab\in I$, either $a\in I$ or some power of $b$ is in $I$. Equivalently, $I$ is primary if every zero-divisor of $R/I$ is nilpotent.
  5. (v)
    Let $I$ be an ideal in $R$. Then $I$ can always be written as a finite union of primary ideals: $I=P_{1}\cap\cdots\cap P_{n}$. Such a finite union is called a primary decomposition of $I$. Given a finite set of generators of $I$, a primary decomposition of $I$ can be effectively computed [Rut92].
  6. (vi)
    Let $S\subset R$ be a set of non zero-divisors of $R$ and denote by $\widetilde{S}\coloneqq\{s_{1}s_{2}\cdots s_{k}\mid k\in\mathbb{N},s_{1},\ldots,s_{k}\in S\}$ its multiplicative closure. The localization $S^{-1}R$ is defined as the ring $\left\{\frac{r}{s}\;\middle|\;r\in R,s\in\widetilde{S}\right\}$, endowed with the operations $\frac{r_{1}}{s_{1}}+\frac{r_{2}}{s_{2}}=\frac{r_{1}s_{2}+r_{2}s_{1}}{s_{1}s_{2}}$ and $\frac{r_{1}}{s_{1}}\cdot\frac{r_{2}}{s_{2}}=\frac{r_{1}r_{2}}{s_{1}s_{2}}$. In particular, $S^{-1}R$ contains $R$ as a subring $\left\{\frac{r}{1}\;\middle|\;r\in R\right\}$. For example, the localization $\{2\}^{-1}\mathbb{Z}$ is the ring of rational numbers whose denominators are powers of $2$. Note that for the canonical map $R\rightarrow S^{-1}R$ to be injective, the set $S$ must not contain zero-divisors.
  7. (vii)
    An $R$-module is defined as an abelian group $(\mathcal{M},+)$ along with an operation $\cdot\;\colon R\times\mathcal{M}\rightarrow\mathcal{M}$, satisfying $r\cdot(m+m^{\prime})=r\cdot m+r\cdot m^{\prime}$, $(r+s)\cdot m=r\cdot m+s\cdot m$, $rs\cdot m=r\cdot(s\cdot m)$ and $1\cdot m=m$. For example, an ideal of $R$ is an $R$-submodule of $R$. For any $d\in\mathbb{N}$, $R^{d}$ is an $R$-module by $s\cdot(r_{1},\ldots,r_{d})=(sr_{1},\ldots,sr_{d})$.
  8. (viii)
    Given two $R$-modules $\mathcal{M},\mathcal{M}^{\prime}$ such that $\mathcal{M}\supseteq\mathcal{M}^{\prime}$, define the quotient $\mathcal{M}/\mathcal{M}^{\prime}\coloneqq\{\overline{m}\mid m\in\mathcal{M}\}$ where $\overline{m_{1}}=\overline{m_{2}}$ if and only if $m_{1}-m_{2}\in\mathcal{M}^{\prime}$. This quotient is also an $R$-module. An $R$-module is called finitely presented if it can be written as a quotient $R^{d}/\langle{v_{1},\ldots,v_{k}}\rangle$ for some $d\in\mathbb{N}$ and some $v_{1},\ldots,v_{k}\in R^{d}$. Here, $\langle{v_{1},\ldots,v_{k}}\rangle$ denotes the $R$-submodule generated by the elements $v_{1},\ldots,v_{k}$.
For $T>0$, denote by $\mathbb{Z}_{/T}[X_{1}^{\pm},\ldots,X_{N}^{\pm}]$ the Laurent polynomial ring over $\mathbb{Z}_{/T}$ with $n$ variables: this is the set of polynomials over the variables $X_{1},X_{1}^{-1},\ldots,X_{N},X_{N}^{-1}$, with coefficients in $\mathbb{Z}_{/T}$, such that $X_{i}X_{i}^{-1}=1$ for all $i$. The following deep theorem by Dong and Shafrir is key to proving our main result in case of prime-power characteristics.
Theorem 2.8 ([T]heorem 1.3).
dong2025s] Let $p$ be a prime number and $e$ be a positive integer. Let $\mathcal{M}$ be a finitely presented module over the Laurent polynomial ring $\mathbb{Z}_{/p^{e}}[X_{1}^{\pm},\ldots,X_{N}^{\pm}]$, and let $m_{0},m_{1},\ldots,m_{K}\in\mathcal{M}$. Then the set of solutions $(z_{11},\ldots,z_{KN})\in\mathbb{Z}^{KN}$ to the equation
$X_{1}^{z_{11}}X_{2}^{z_{12}}\cdots X_{N}^{z_{1N}}\cdot m_{1}+\cdots+X_{1}^{z_{K1}}X_{2}^{z_{K2}}\cdots X_{N}^{z_{KN}}\cdot m_{K}=m_{0}$ (2.2)
is effectively $p$-normal.
Finally, we say a few words about the relation between linear recurrence sequences and formal power series. Let $R$ be a commutative ring. A (univariate) formal power series over $R$ is a formal sum $\sum_{n=0}^{\infty}a_{n}Y^{n}=a_{0}+a_{1}Y+a_{2}Y^{2}+\cdots$, where each $a_{n}$ is an element in $R$. The set of all formal power series over $R$ forms a ring and is denoted by $R[[Y]]$.
Given a linear recurrence sequence $\gamma=(\gamma_{0},\gamma_{1},\gamma_{2},\ldots)\in R^{\mathbb{N}}$ satisfying the recursion
$\forall n\geq d,\quad\gamma_{n}=a_{1}\gamma_{n-1}+\cdots+a_{d}\gamma_{n-d},$
the characteristic polynomial of $\gamma$ is defined as
$f(Y)\coloneqq Y^{d}-a_{1}Y^{d-1}-\cdots-a_{d-1}Y-a_{d}\in R[Y].$
Consider the polynomial
$\phi(Y)\coloneqq Y^{d}\cdot f\left(\frac{1}{Y}\right)=1-a_{1}Y-\cdots-a_{d-1}Y^{d-1}-a_{d}Y^{d},$
and the formal power series
$g(Y)\coloneqq\sum_{i=0}^{\infty}\gamma_{i}Y^{i}\in R[[Y]].$
The product
$\phi(Y)g(Y)=\sum_{i=0}^{\infty}\gamma_{i}Y^{i}-a_{1}\sum_{i=1}^{\infty}\gamma_{i-1}Y^{i}-\cdots-a_{d}\sum_{i=d}^{\infty}\gamma_{i-d}Y^{i}$
is equal to a polynomial $h(Y)\in R[Y]$, since all its coefficients of degree $\geq d$ is zero. Therefore $g(Y)=\frac{h(Y)}{\phi(Y)}$.
3 Skolem Problem in rings of positive characteristic
In this section we prove Theorem 1.1. For a linear recurrence sequence $\gamma=(\gamma_{0},\gamma_{1},\gamma_{2},\ldots)\in\mathcal{R}^{\mathbb{N}}$, denote by $\mathfrak{Z}(\gamma)$ its zero set $\{n\in\mathbb{N}\mid\gamma_{n}=0\}$. We now fix a ring $\mathcal{R}$ of characteristic $T>0$ and a linear recurrence sequence $\gamma$, and Theorem 1.1 boils down to deciding whether $\mathfrak{Z}(\gamma)$ is empty.
Let $T=p_{1}^{e_{1}}\cdots p_{k}^{e_{k}}$ be the prime factorization of $T$. Since $p_{1}^{e_{1}}\mathcal{R},\ldots,p_{k}^{e_{k}}\mathcal{R}$ are pairwise coprime ideals of $\mathcal{R}$ (i.e., $1\in p_{i}^{e_{i}}\mathcal{R}+p_{j}^{e_{j}}\mathcal{R}$ for $i\neq j$), we have
$p_{1}^{e_{1}}\mathcal{R}\cap\cdots\cap p_{k}^{e_{k}}\mathcal{R}=\{0\}$ (3.1)
(c.f. [Mat89, Theorem 1.3]). For each $i=1,\ldots,k$, consider the quotient $\mathcal{A}_{i}\coloneqq\mathcal{R}/p_{i}^{e_{i}}\mathcal{R}$ and the projection map $\rho_{i}\colon\mathcal{R}\rightarrow\mathcal{R}/p_{i}^{e_{i}}\mathcal{R}$. Then $\alpha_{i}\coloneqq\rho_{i}(\gamma)$ is a linear recurrence sequence over the ring $\mathcal{A}_{i}$. By Equation (3.1), we have $r=0$ if and only if $\rho_{i}(r)=0$ for all $i=1,\ldots,k$. Therefore
$\mathfrak{Z}(\gamma)=\mathfrak{Z}(\alpha_{1})\cap\cdots\cap\mathfrak{Z}(\alpha_{k}).$
We will first show that each $\mathfrak{Z}(\alpha_{i})$ is effectively $p_{i}$-normal for $i=1,\ldots,k$, and then show that their intersection is effectively computable.
3.1 Effective $p_{i}$-normality of $\mathfrak{Z}(\alpha_{i})$
Fix an index $i\in\{1,\ldots,k\}$. For brevity, we now omit the subscript $i$, so $\alpha$ is a linear recurrence sequence over $\mathcal{A}$, where $p^{e}\mathcal{A}=0$ for a prime number $p$ and positive integer $e$. In this subsection we will prove: See 1.2
Let
$f(X)\coloneqq X^{d}-a_{1}X^{d-1}-\cdots-a_{d-1}X-a_{d}\in\mathcal{A}[X]$
be the characteristic polynomial of $\alpha$. Observe that by extending the ring $\mathcal{A}$, one can without loss of generality suppose $f$ splits in $\mathcal{A}$:
Lemma 3.1 (folklore).
One can effectively compute a ring $\widetilde{\mathcal{A}}\supseteq\mathcal{A}$, finitely generated as an $\mathcal{A}$-module, such that $f(X)=\prod_{i=1}^{t}(X-r_{i})^{d_{i}}$ with $t\in\mathbb{N},r_{1},\ldots,r_{t}\in\widetilde{\mathcal{A}}$.
Proof.
Let $T$ be a new variable and consider the quotient ring $\mathcal{A}[T]/\langle{f(T)}\rangle$, where $\langle{f(T)}\rangle$ denotes the ideal generated by $f(T)$. The ring $\mathcal{A}[T]/\langle{f(T)}\rangle$ is finitely generated as an $\mathcal{A}$-module by $\{1,T,\ldots,T^{d-1}\}$. Since $f$ is monic, the map $\mathcal{A}\rightarrow\mathcal{A}[T]/\langle{f(T)}\rangle$ is injective. Over the base ring $\mathcal{A}[T]/\langle{f(T)}\rangle$, the polynomial $f(X)$ has a linear factor, namely XTX-T. We can then replace $f(X)$ by $\frac{f(X)}{X-T}$ and obtain the lemma by induction on the degree of $f$. ∎
From now on we suppose that $f$ splits as a product $\prod_{i=1}^{t}(X-r_{i})^{d_{i}}$ over $\mathcal{A}$. When $\mathcal{A}$ is a field or an integral domain, it is a standard result [OW15] that the $n$-th term of the sequence $\alpha$ can be written as an exponential polynomial $\alpha_{n}=\sum_{i=1}^{t}r_{i}^{n}g_{i}(n)$ with $g_{i}(X)\in\mathcal{A}[X]$, (for example by considering the Jordan normal form of the transition matrix of $\alpha$). However this no longer holds in the presence of zero-divisors in the ring $\mathcal{A}$, since these zero-divisors can neither be inverted nor localized. To overcome this, we will use primary decomposition to reduce to the case where all zero-divisors are nilpotent.
Let $\langle{0}\rangle=P_{1}\cap\cdots\cap P_{s}$ be the primary decomposition of the zero ideal $\langle{0}\rangle\subseteq\mathcal{A}$. Consider the projection maps $\tau_{j}\colon\mathcal{A}\rightarrow\mathcal{A}/P_{j}$. Since $\langle{0}\rangle=P_{1}\cap\cdots\cap P_{s}$, an element $a\in\mathcal{A}$ is zero if and only if $\tau_{j}(a)=0$ for all $j=1,\ldots,s$. Therefore
$\mathfrak{Z}(\alpha)=\mathfrak{Z}(\tau_{1}(\alpha))\cap\cdots\cap\mathfrak{Z}(\tau_{s}(\alpha)).$
Since the intersection of $p$-normal subsets is effectively $p$-normal (Lemma 2.4), to show $\mathfrak{Z}(\alpha)$ is $p$-normal it suffices to show that each $\mathfrak{Z}(\tau_{j}(\alpha)),j=1,\ldots,s$ is $p$-normal. Thus, we are reduced to considering the case where $\mathcal{A}$ is of the form $A/P$, where $A$ is a ring with $p^{e}A=0$, and $P$ is a primary ideal of $A$. In particular, every zero-divisor of $\mathcal{A}=A/P$ is nilpotent (Definition 2.7(iv)).
Lemma 3.2.
Suppose every zero-divisor of $\mathcal{A}$ is nilpotent, and the characteristic polynomial $f$ of $(\alpha_{n})_{n\in\mathbb{N}}$ splits as a product $\prod_{i=1}^{t}(X-r_{i})^{d_{i}}$ with $t\in\mathbb{N},r_{1},\ldots,r_{t}\in\mathcal{A}$. Then there is a finite, effectively computable set of non zero-divisors $S\subseteq\mathcal{A}$, numbers $u,N\in\mathbb{N}$, a subset $\mathcal{I}\subseteq\{1,\ldots,t\}$, and $c_{ij}\in S^{-1}\mathcal{A},i\in\mathcal{I};j=1,\ldots,u$, such that
$\alpha_{n}=\sum_{i\in\mathcal{I}}r_{i}^{n}\sum_{j=0}^{u}c_{ij}n^{j}$ (3.2)
for all $n\geq N$. Furthermore, every $r_{i},i\in\mathcal{I}$ is invertible in $S^{-1}\mathcal{A}$.
Proof.
Let $\mathcal{I}$ be the set $\{i\mid r_{i}\text{ is not a zero-divisor}\}$, and let
$S=\{r_{i}-r_{j}\mid r_{i}-r_{j}\text{ is not a zero-divisor}\}\cup\{r_{i}\mid i\in\mathcal{I}\}.$
Then every $r_{i},i\in\mathcal{I}$ is invertible in $S^{-1}\mathcal{A}$. Denote $\widetilde{\mathcal{A}}\coloneqq S^{-1}\mathcal{A}$. Consider the formal power series ring $\widetilde{\mathcal{A}}[[Y]]$ and let $g(Y)\coloneqq\sum_{i=0}^{\infty}\gamma_{i}Y^{i}$. Note that elements of $\widetilde{\mathcal{A}}[[Y]]$ whose constant term is one are invertible. Denote
$\phi(Y)\coloneqq Y^{d}\cdot f\left(\frac{1}{Y}\right)=\prod_{i=1}^{t}(1-r_{i}Y)^{d_{i}}=1-a_{1}Y-\cdots-a_{d-1}Y^{d-1}-a_{d}Y^{d}.$
Then $g(Y)=\frac{h(Y)}{\phi(Y)}$ for some polynomial $h(Y)\in\widetilde{\mathcal{A}}[Y]$ (see Section 2).
We claim that $\frac{1}{\phi(Y)}=\frac{1}{\prod_{i=1}^{t}(1-r_{i}Y)^{d_{i}}}$ is effectively equal to a sum
$\sum_{i=1}^{t}\sum_{j=1}^{u}a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}$ (3.3)
for some $u\in\mathbb{N}$, $a_{ij}\in\widetilde{\mathcal{A}}[Y],1\leq i\leq t,0\leq j\leq u$. The proof is done in two steps.
Step 1: we reduce to the case where all $r_{i}-r_{j},1\leq i<j\leq t$ are zero-divisors. Suppose there are $r_{i},r_{j}$ such that $r_{i}-r_{j}$ is not a zero-divisor. Without loss of generality suppose these are $r_{1},r_{2}$. Then $r_{1}-r_{2}\in S$, so $r_{1}-r_{2}$ is invertible in $\widetilde{\mathcal{A}}=S^{-1}\mathcal{A}$. We have
$\frac{1}{(1-r_{1}Y)(1-r_{2}Y)}=\frac{r_{1}}{r_{1}-r_{2}}\cdot\frac{1}{1-r_{1}Y}+\frac{-r_{2}}{r_{1}-r_{2}}\cdot\frac{1}{1-r_{2}Y}.$
Divide both sides by $(1-r_{1}Y)^{d_{1}-1}(1-r_{2}Y)^{d_{2}-1}\prod_{i=3}^{t}(1-r_{i}Y)^{d_{i}}$, we obtain that $\frac{1}{\phi(Y)}=\frac{1}{\prod_{i=1}^{t}(1-r_{i}Y)^{d_{i}}}$ is equal to a sum of two fractions
$\frac{r_{1}}{r_{1}-r_{2}}\cdot\frac{1}{(1-r_{2}Y)^{d_{2}-1}\prod_{i\neq 2}(1-r_{i}Y)^{d_{i}}}+\frac{-r_{2}}{r_{1}-r_{2}}\cdot\frac{1}{(1-r_{1}Y)^{d_{1}-1}\prod_{i\neq 1}(1-r_{i}Y)^{d_{i}}}.$ (3.4)
We can repeat this argument for each of the two fractions in (3.4). Note that the two denominators in (3.4) have smaller degree than $\prod_{i=1}^{t}(1-r_{i}Y)^{d_{i}}$. Therefore, we can apply this argument iteratively until each denominator is of the form $\prod_{i\in J}(1-r_{i}Y)^{d^{\prime}_{i}}$ for some $J\subseteq\{1,\ldots,t\}$, where all $r_{i}-r_{j},i,j\in J$ are zero-divisors. This reduces to the next case.
Step 2: we now consider the case where all $r_{i}-r_{j},1\leq i<j\leq t$ are zero-divisors. We will show that $\frac{1}{\phi(Y)}$ can be written in the form (3.3), using induction on $t$. The base case $t=1$ is trivial. For the induction step, suppose $t\geq 2$. Since $r_{1}-r_{2}$ is a zero-divisor and hence nilpotent, there exists $\ell\in\mathbb{N}$ such that $(r_{1}-r_{2})^{\ell}=0$. Then
$\frac{1}{(1-r_{1}Y)(1-r_{2}Y)}$ $=\frac{1}{(1-r_{2}Y)^{2}}\cdot\frac{1}{\frac{1-r_{1}Y}{1-r_{2}Y}}$
$=\frac{1}{(1-r_{2}Y)^{2}}\cdot\frac{1}{1-\frac{(r_{1}-r_{2})Y}{1-r_{2}Y}}$
$=\frac{1}{(1-r_{2}Y)^{2}}\cdot\left(1+\frac{(r_{1}-r_{2})Y}{1-r_{2}Y}+\cdots+\frac{(r_{1}-r_{2})^{\ell-1}Y^{\ell-1}}{(1-r_{2}Y)^{\ell-1}}\right)$
$=\frac{1}{(1-r_{2}Y)^{2}}+\frac{(r_{1}-r_{2})Y}{(1-r_{2}Y)^{3}}+\cdots+\frac{(r_{1}-r_{2})^{\ell-1}Y^{\ell-1}}{(1-r_{2}Y)^{\ell+1}}.$
Divide both sides by $(1-r_{1}Y)^{d_{1}-1}(1-r_{2}Y)^{d_{2}-1}\prod_{i=3}^{t}(1-r_{i}Y)^{d_{i}}$, we obtain that $\frac{1}{\phi(Y)}=\frac{1}{\prod_{i=1}^{t}(1-r_{i}Y)^{d_{i}}}$ is equal to a sum of $\ell$ fractions
$\frac{1}{(1-r_{1}Y)^{d_{1}-1}(1-r_{2}Y)^{d_{2}+1}\prod_{i=3}^{t}(1-r_{i}Y)^{d_{i}}}+\cdots+\frac{(r_{1}-r_{2})^{\ell-1}Y^{\ell-1}}{(1-r_{1}Y)^{d_{1}-1}(1-r_{2}Y)^{d_{2}+\ell}\prod_{i=3}^{t}(1-r_{i}Y)^{d_{i}}}.$ (3.5)
We can repeat this argument for each of the $\ell$ fractions in (3.5). Note that the multiplicity of the factor $1-r_{1}Y$ in the denominator of each fraction is smaller than $d_{1}$. Therefore, we can apply this argument iteratively until this multiplicity becomes zero. This eliminates the factor $1-r_{1}Y$, thereby decreasing $t$ by one. We thereby conclude the induction step.
Combining the two steps, we have obtained
$\frac{1}{\phi(Y)}=\sum_{i=1}^{t}\sum_{j=1}^{u}a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}.$
Consider a term $a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}$, we have
$\frac{1}{(1-r_{i}Y)^{j}}=\sum_{n=0}^{\infty}\binom{j+n-1}{j-1}r_{i}^{n}Y^{n}.$
Write $a_{ij}=\sum_{s=0}^{D_{ij}}b_{s}Y^{s}$. If $r_{i}$ is not a zero-divisor, it is invertible in $\widetilde{\mathcal{A}}=S^{-1}\mathcal{A}$, so
$a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}=\sum_{n=0}^{\infty}\sum_{s=0}^{D_{ij}}\binom{j+n-s-1}{j-1}(b_{s}r_{i}^{-s})r_{i}^{n}Y^{n}$
and its $n$-th coefficient is some exponential-polynomial sum $r_{i}^{n}\sum_{j=0}^{u}c_{ij}n^{j}$. If $r_{i}$ is a zero-divisor and hence nilpotent, then $r_{i}^{\ell_{i}}=0$ for some effectively computable $\ell_{i}\in\mathbb{N}$, so $a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}$ is a polynomial of degree at most $\ell_{i}+D_{ij}$. This allows us to conclude the proof of the lemma by taking $N\coloneqq 1+\max_{i\not\in\mathcal{I},0\leq j\leq u}\{\ell_{i}+D_{ij}\}$. ∎
From now on, we can without loss of generality replace $\mathcal{A}$ with its extension $\widetilde{\mathcal{A}}=S^{-1}\mathcal{A}$ from in Lemma 3.2, and write
$\alpha_{n}=\sum_{i\in\mathcal{I}}r_{i}^{n}\sum_{j=0}^{u}c_{ij}n^{j}$
for all $n\geq N$, with $c_{ij},r_{i}\in\mathcal{A}$ for all $i,j$. In particular, each $r_{i}$ is invertible.
A linear recurrence $\alpha$ is called simple if $u=0$ in the above expression. Since $p^{e}=0$ in $\mathcal{A}$, we have $n^{j}=(n+p^{e})^{j}$ in $\mathcal{A}$. Therefore the linear recurrence sequence $\alpha$ can be written as a union of $p^{e}$ simple recurrences $\left(\alpha_{p^{e}n}\right)_{n\in\mathbb{N}},\left(\alpha_{p^{e}n+1}\right)_{n\in\mathbb{N}},\ldots,\left(\alpha_{p^{e}n+(p^{e}-1)}\right)_{n\in\mathbb{N}}$. In particular, we have
$\alpha_{p^{e}n+q}=\sum_{i\in\mathcal{I}}r_{i}^{p^{e}n+q}\sum_{j=0}^{u}c_{ij}(p^{e}n+q)^{j}=\sum_{i\in\mathcal{I}}\left(r_{i}^{p^{e}}\right)^{n}d_{i,q},$
with $d_{i,q}\coloneqq r_{i}^{q}\sum_{j=0}^{u}c_{ij}q^{j}$. Therefore we can without loss of generality replace each $r_{i}$ with $r_{i}^{p^{e}}$ and consider each simple recurrence.
Let
$\alpha_{n}=\sum_{i\in\mathcal{I}}c_{i}r_{i}^{n}$ (3.6)
be a simple recurrence sequence. Since each $r_{i}$ is invertible, we can extend the definition of the sequence (3.6) from $n\geq N$ to all $n\in\mathbb{Z}$. We now show that the solution set of $\sum_{i\in\mathcal{I}}c_{i}r_{i}^{z}=0$ is effectively $p$-normal in $\mathbb{Z}$.
Lemma 3.3.
Let $\mathcal{A}$ be a ring where $p^{e}=0$ and $\mathcal{I}$ be a finite set of indices. Let $c_{i},r_{i}\in\mathcal{A}$ for all $i\in\mathcal{I}$, such that each $r_{i}$ is invertible. Then, the set of integers $z\in\mathbb{Z}$ such that $\sum_{i\in\mathcal{I}}c_{i}r_{i}^{z}=0$, is an effectively $p$-normal subset of $\mathbb{Z}$.
Proof.
Write $\mathcal{I}$ as $\{1,\ldots,t\}$ for some $t\in\mathbb{N}$. We can give $\mathcal{A}$ a $\mathbb{Z}[X_{1}^{\pm},\ldots,X_{t}^{\pm}]$-module structure by letting $X_{i}$ act as $r_{i}$. Denote by $\mathcal{M}$ the submodule generated by $c_{1},\ldots,c_{t}$, we have $p^{e}\mathcal{M}=0$. Using the standard algorithm in commutative algebra (see [BCMI81, Theorem 2.14] or [BCR94, Theorem 2.6]) we can compute a finite presentation of $\mathcal{M}$ as a $\mathbb{Z}[X_{1}^{\pm},\ldots,X_{t}^{\pm}]$-module. Then, the equation $\sum_{i=1}^{t}c_{i}r_{i}^{z}=0$ in $\mathcal{A}$ is equivalent to the equation $\sum_{i=1}^{t}X_{i}^{z}\cdot c_{i}=0$ in the $\mathbb{Z}[X_{1}^{\pm},\ldots,X_{t}^{\pm}]$-module $\mathcal{M}$.
Consider the set of solutions $(z_{11},z_{12},\ldots,z_{tt})\in\mathbb{Z}^{t^{2}}$ to the S-unit equation
$X_{1}^{z_{11}}X_{2}^{z_{12}}\cdots X_{t}^{z_{1t}}\cdot c_{1}+\cdots+X_{1}^{z_{t1}}X_{2}^{z_{t2}}\cdots X_{t}^{z_{tt}}\cdot c_{t}=0,$
in $\mathcal{M}$. By Theorem 2.8, the solution set $S$ is effectively $p$-normal in $\mathbb{Z}^{t^{2}}$. Let $\boldsymbol{b}=(b_{11},b_{12},\ldots,b_{tt})\in\mathbb{Z}^{t^{2}}$ be the vector defined by $b_{ij}=1$ for $i=j$, and $b_{ij}=0$ for $i\neq j$. Applying Corollary 2.6 to the solution set $S$ and the vector $\boldsymbol{b}$, we obtain that the set of integers $z$ satisfying
$X_{1}^{z}\cdot c_{1}+\cdots+X_{t}^{z}\cdot c_{t}=0,$
is effectively $p$-normal in $\mathbb{Z}$. This concludes the proof. ∎
This allows us to conclude the proof of Proposition 1.2:
Proof of Proposition 1.2.
The previous discussion showed that we can reduce to the case where $\alpha_{n}=\sum_{i=1}^{t}c_{i}r_{i}^{n}$ for all $n\geq N$, where each $r_{i}$ is invertible. For each $0\leq n<N$, we can verify whether $\alpha_{n}=0$ individually. Lemma 3.3 showed that the zero set
$S=\left\{z\in\mathbb{Z}\;\middle|\;\sum_{i=1}^{t}c_{i}r_{i}^{z}=0\right\}$
is effectively $p$-normal in $\mathbb{Z}$. Since $\mathfrak{Z}(\alpha)=\left(S\cap[N,\infty)\right)\cup F$, where $F=\{0\leq n<N\mid\alpha_{n}=0\}$, we conclude that $\mathfrak{Z}(\alpha)$ is an effectively $p$-normal subset of $\mathbb{N}$. ∎
3.2 Intersecting $p_{i}$-normal sets
In this subsection we prove the following proposition, which will yield Theorem 1.1 when combined with Proposition 1.2. See 1.3
We start with a deep result of Karimov, Luca, Nieuwveld, Ouaknine, and Worrell:
Lemma 3.4 ([KLN+25, Theorem 3.2]).
Let $p,q\in\mathbb{Z}_{>0}$ be multiplicatively independent, and let $a_{1},\ldots,a_{k},b_{1},\ldots,b_{m},d\in\mathbb{Z}$. Denote by $\mathcal{S}$ the set of all $(n_{1},\ldots,n_{k},n^{\prime}_{1},\ldots,n^{\prime}_{m})\in\mathbb{N}^{k+m}$ to the equation
$p^{n_{1}}a_{1}+\cdots+p^{n_{k}}a_{k}+q^{n^{\prime}_{1}}b_{1}+\cdots+q^{n^{\prime}_{m}}b_{m}=d,$ (3.7)
Then the set $\mathcal{S}$ is of the form
$\bigcup_{i\in I}\bigcap_{j\in J_{i}}\mathcal{X}_{j},$ (3.8)
where $I$ and each $J_{i},i\in I$ is finite, and each set $\mathcal{X}_{j}$ is in of one of the following forms
  1. (i)
    $\mathcal{X}_{j}=\{(n_{1},\ldots,n^{\prime}_{m})\in\mathbb{N}^{k+m}\mid n_{s}=n_{t}+c\}$,
  2. (ii)
    $\mathcal{X}_{j}=\{(n_{1},\ldots,n^{\prime}_{m})\in\mathbb{N}^{k+m}\mid n^{\prime}_{s}=n^{\prime}_{t}+c\}$,
  3. (iii)
    $\mathcal{X}_{j}=\{(n_{1},\ldots,n^{\prime}_{m})\in\mathbb{N}^{k+m}\mid n_{s}=c\}$,
  4. (iv)
    $\mathcal{X}_{j}=\{(n_{1},\ldots,n^{\prime}_{m})\in\mathbb{N}^{k+m}\mid n^{\prime}_{s}=c\}$,
with integer $c$. Furthermore, a representation of $\mathcal{S}$ in the form (3.7) can be effectively computed.
From Lemma 3.4, we obtain the following:
Lemma 3.5.
Let $p,q\in\mathbb{Z}_{>0}$ be multiplicatively independent integers, and let $a_{1},\ldots,a_{k}$, $b_{1},\ldots,b_{m}$, $d\in\mathbb{Q}$. Then, there exists an effectively computable number $C\in\mathbb{N}$, such that all solutions $(n_{1},\ldots,n_{k},n^{\prime}_{1},\ldots,n^{\prime}_{m})\in\mathbb{N}^{k+m}$ to the equation
$p^{n_{1}}a_{1}+\cdots+p^{n_{k}}a_{k}+q^{n^{\prime}_{1}}b_{1}+\cdots+q^{n^{\prime}_{m}}b_{m}=d,$ (3.9)
satisfy
$\left|p^{n_{1}}a_{1}+\cdots+p^{n_{k}}a_{k}\right|\leq C.$ (3.10)
Proof.
By multiplying $a_{1},\ldots,a_{k}$, $b_{1},\ldots,b_{m}$, $d$ with their common denominator, we can without loss of generality suppose them to be integers. We use induction on $k+m$. When $k=0$ or $m=0$, the bound (3.10) holds for $C=d$. For the induction step, suppose $k\geq 1$ and $m\geq 1$. By Lemma 3.4, we can effectively represent the solution set of Equation (3.9) in the form (3.8). For each $i\in I$, let $S_{i}\coloneqq\bigcap_{j\in J_{i}}\mathcal{X}_{j}$. It suffices to compute a bound $C_{i}$ for each $i\in I$, such that $\left|p^{n_{1}}a_{1}+\cdots+p^{n_{k}}a_{k}\right|\leq C_{i}$ for all $(n_{1},\ldots,n_{m}^{\prime})\in S_{i}$. Fix an index $i\in I$. If $S_{i}=\mathbb{N}^{k+m}$ then $a_{1},\ldots,a_{k}$, $b_{1},\ldots,b_{m}$, $d$ must all be zero, and the lemma becomes trivial. Therefore suppose $S_{i}\neq\mathbb{N}^{k+m}$. Take any $\mathcal{X}_{j},j\in J_{i}$, then $\mathcal{X}_{j}$ is in of one of the four forms in Lemma 3.4. In case (i), every solution in $S_{i}=\bigcap_{j\in J_{i}}\mathcal{X}_{j}$ satisfies $n_{s}=n_{t}+c$, so we can replace $n_{s}$ with $n_{t}+c$ in Equation (3.9). Thereby, we have replaced $p^{n_{s}}a_{s}+p^{n_{t}}a_{t}$ with $p^{n_{t}}(p^{c}a_{s}+a_{t})$ in Equation (3.9), thereby eliminating the variable $n_{s}$ and decreasing $k$ by one. By the induction hypothesis, we can compute an effective bound (3.10) for the solutions of this new equation. In cases (ii), (iii), (iv), a bound can be obtained similarly. ∎
Lemma 3.5 yields the special case of Proposition 1.3 with $k=2$:
Lemma 3.6.
Let $p_{1},p_{2}$ be multiplicatively independent. Let $S_{1},S_{2}$ be respectively $p_{1}$-normal and $p_{2}$-normal subsets of $\mathbb{N}$. Then $S_{1}\cap S_{2}$ is effectively equal to a union $T_{1}\cup T_{2}$, where $T_{1}$ and $T_{2}$ are respectively $p_{1}$-normal and $p_{2}$-normal subsets of $\mathbb{N}$.
Proof.
For $i=1,2$, we have $S_{i}=\left(S^{\prime}_{i}\cap[n_{i},\infty)\right)\cup F_{i}$ where $S_{1}^{\prime},S_{2}^{\prime}$ are respectively $p_{1}$-normal and $p_{2}$-normal subsets of $\mathbb{Z}$, and $F_{1}\subseteq\{0,1,\ldots,n_{1}-1\},F_{2}\subseteq\{0,1,\ldots,n_{2}-1\}$. By enlarging $F_{1},F_{2}$ we can without loss of generality suppose $S_{i}=\left(S^{\prime}_{i}\cap[n,\infty)\right)\cup F_{i},i=1,2,$ with $n=\max\{n_{1},n_{2}\}$ and $F_{1},F_{2}\subseteq\{0,1,\ldots,n-1\}$. We now show that $S^{\prime}_{1}\cap S^{\prime}_{2}$ is effectively a union of $p_{1}$-normal and $p_{2}$-normal subsets of $\mathbb{Z}$.
By Observation 2.2, each $S^{\prime}_{i}$ is a finite union of elementary $p_{i}$-nested sets of $\mathbb{Z}$ and sets of the form $a\mathbb{Z}+b$. Therefore it suffices to consider the intersection $U_{1}\cap U_{2}$, where $U_{i},i=1,2$, is elementary $p_{i}$-nested or $a\mathbb{Z}+b$. We show that $U_{1}\cap U_{2}$ is either effectively $p_{1}$-normal or effectively $p_{2}$-normal. There are four cases.
Case 1: $U_{1}$ is elementary $p_{1}$-nested and $U_{2}$ is elementary $p_{2}$-nested. Write
$U_{1}=\{a_{0}+p_{1}^{\ell_{1}k_{1}}a_{1}+\cdots+p_{1}^{\ell_{1}k_{r}}a_{r}\mid k_{1},k_{2},\ldots,k_{r}\in\mathbb{N}\},$
and
$U_{2}=\{b_{0}+p_{2}^{\ell_{2}k^{\prime}_{1}}b_{1}+\cdots+p_{2}^{\ell_{2}k^{\prime}_{s}}b_{s}\mid k^{\prime}_{1},k^{\prime}_{2},\ldots,k^{\prime}_{s}\in\mathbb{N}\}.$
Take $p\coloneqq p_{1}^{\ell_{1}}$ and $q\coloneqq p_{2}^{\ell_{2}}$, they are multiplicatively independent. Any $z\in U_{1}\cap U_{2}$ satisfies
$z=a_{0}+p^{k_{1}}a_{1}+\cdots+p^{k_{r}}a_{r}=b_{0}+q^{k^{\prime}_{1}}b_{1}+\cdots+q^{k^{\prime}_{s}}b_{s}.$
Therefore by Lemma 3.5, we can effectively compute a bound $C\in\mathbb{N}$ such that every $z\in U_{1}\cap U_{2}$ satisfies $|z|\leq C$. Therefore, $U_{1}\cap U_{2}$ is effectively bounded (so it is a finite union of singletons), and is therefore both effectively $p_{1}$-normal and $p_{2}$-normal.
Case 2: $U_{1}$ is elementary $p_{1}$-nested and $U_{2}$ is $a\mathbb{Z}+b$. In this case, both $U_{1}$ and $U_{2}$ are $p_{1}$-normal, therefore their intersection is effectively $p_{1}$-normal (Lemma 2.4).
Case 3: $U_{1}$ is $a\mathbb{Z}+b$ and $U_{2}$ elementary $p_{2}$-nested. Similar to the previous case, $U_{1}\cap U_{2}$ is effectively $p_{2}$-normal.
Case 4: $U_{1}=a\mathbb{Z}+b$ and $U_{2}=a^{\prime}\mathbb{Z}+b^{\prime}$. In this case $U_{1}\cap U_{2}$ is either empty or of the form $A\mathbb{Z}+B$, where $A$ is the least common multiplier of $a$ and $a^{\prime}$. Therefore $U_{1}\cap U_{2}$ is both $p_{1}$-normal and $p_{2}$-normal.
In all four cases, the intersection $U_{1}\cap U_{2}$ is either effectively $p_{1}$-normal or effectively $p_{2}$-normal. Note that a finite union of $p$-normal sets is $p$-normal. Since $S^{\prime}_{1}\cap S^{\prime}_{2}$ is a finite union of sets of the form $U_{1}\cap U_{2}$, it is effectively equal to a union $T_{1}^{\prime}\cup T_{2}^{\prime}$, where $T_{1}^{\prime}\subseteq\mathbb{Z}$ is effectively $p_{1}$-normal and $T_{2}^{\prime}\subseteq\mathbb{Z}$ is effectively $p_{2}$-normal. Since
$S_{1}\cap S_{2}=\left(S^{\prime}_{1}\cap S^{\prime}_{2}\cap[n,\infty)\right)\cup(F_{1}\cap F_{2})=\left(T_{1}^{\prime}\cap[n,\infty)\right)\cup\left(T_{2}^{\prime}\cap[n,\infty)\right)\cup(F_{1}\cap F_{2}),$
we conclude that $S_{1}\cap S_{2}$ is effectively a union of the $p_{1}$-normal subset $T_{1}\coloneqq T_{1}^{\prime}\cap[n,\infty)\subseteq\mathbb{N}$ and the $p_{2}$-normal subset $T_{2}\coloneqq\left(T_{2}^{\prime}\cap[n,\infty)\right)\cup(F_{1}\cap F_{2})\subseteq\mathbb{N}$. ∎
Lemma 3.6 allows us to prove Proposition 1.3 in full generality:
Proof of Proposition 1.3.
We use induction on $k$. The base case $k=2$ follows directly from Lemma 3.6. For the induction step, for each $i=1,\ldots,k$, let $S_{i}$ be a $p_{i}$-normal subset of $\mathbb{N}$. By Lemma 3.6, we can effectively write $S_{1}\cap S_{2}=T_{1}\cup T_{2}$, where $T_{1},T_{2}$ are respectively $p_{1}$-normal and $p_{2}$-normal sets. Then
$S_{1}\cap S_{2}\cap S_{3}\cap\cdots\cap S_{k}=(T_{1}\cap S_{3}\cap\cdots\cap S_{k})\cup(T_{2}\cap S_{3}\cap\cdots\cap S_{k}).$
By the induction hypothesis on $k$, both $T_{1}\cap S_{3}\cap\cdots\cap S_{k}$ and $T_{2}\cap S_{3}\cap\cdots\cap S_{k}$ can be effectively written as a union of $p_{i}$-normal sets with $i=1,\ldots,k$. Therefore $S_{1}\cap S_{2}\cap S_{3}\cap\cdots\cap S_{k}$ is also effectively a union $T_{1}\cup\cdots\cup T_{k}$ of $p_{i}$-normal sets with $i=1,\ldots,k$. In particular, it is empty if and only if each $T_{i}$ is empty. ∎
Combining Proposition 1.2 and 1.3, we immediately obtain Theorem 1.1:
Proof of Theorem 1.1.
Let $T=p_{1}^{e_{1}}\cdots p_{k}^{e_{k}}$ be the prime factorization of $T$, then $\mathfrak{Z}(\gamma)=\mathfrak{Z}(\alpha_{1})\cap\cdots\cap\mathfrak{Z}(\alpha_{k}),$ where $\alpha_{i}$ is the linear recurrence sequence $(\gamma\mod p_{i}^{e_{i}})$ over the quotient ring $\mathcal{R}/p_{i}^{e_{i}}\mathcal{R}$. By Proposition 1.2, the zero set $\mathfrak{Z}(\alpha_{i})$ is effectively $p_{i}$-normal. Then by Proposition 1.3, the intersection $\mathfrak{Z}(\alpha_{1})\cap\cdots\cap\mathfrak{Z}(\alpha_{k})$ is effectively equal to a union of $p_{i}$-normal subsets of $\mathbb{N}$ with $i=1,\ldots,k$, whose emptiness can be tested. This yields a procedure to test whether $\mathfrak{Z}(\gamma)$ is empty. ∎
It is worth noting that the proof of Theorem 1.1 also shows that the zero set $\mathfrak{Z}(\gamma)$ is effectively a union of $p_{i}$-normal sets with $i=1,\ldots,k$.